OpenStudy (anonymous):

Evaluate f^-2g for f = 3 and g = 27. a) 1/3 b) 3 c) 9

OpenStudy (anonymous):

\[f ^{-2}g = ?\] We've looked at negative exponents in two other problems. Can you tell me how this would simplify?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

If it is a negative exponent then that means it goes in the denominator and is raised to the power so if f = 3 then: \[f^{-2} = 1/(f^{2})\] = 1/9 if g = 27 then 27/9 = 3